Cupcake123456:

Noah started to solve the equation –4.6p – 6.3p + 3.9 = –9.18 below. Combine like terms: –10.9p + 3.9 = –9.18 Apply the next steps to solve the equation. What is the solution? p =

Cupcake123456:

would it b 1.2?

GoodFellow:

"combine like terms" yes you are on the right track. believe in yourself

jhonyy9:

what is the next step ?

jhonyy9:

you need isolate the term of p on the left side so what make for this ?

Cupcake123456:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 what is the next step ? \(\color{#0cbb34}{\text{End of Quote}}\) - 10.9 p + 3.9 = - 9.18

Cupcake123456:

? i think

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Cupcake123456 \(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 what is the next step ? \(\color{#0cbb34}{\text{End of Quote}}\) - 10.9 p + 3.9 = - 9.18 \(\color{#0cbb34}{\text{End of Quote}}\) but what will be with this 3,9 on the left side ?

jhonyy9:

so to eliminate it subtract from both sides 3,9 and so what will get ?

Cupcake123456:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 so to eliminate it subtract from both sides 3,9 and so what will get ? \(\color{#0cbb34}{\text{End of Quote}}\) 10.9 p = - 9.18 - 3.9?

jhonyy9:

yes continue pls.

jhonyy9:

you need solve it for p

Cupcake123456:

10.9 p = - 13.08?

Cupcake123456:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 you need solve it for p \(\color{#0cbb34}{\text{End of Quote}}\) 10.9 p = - 13.08?

jhonyy9:

ok and now how you get the value of p ?